OpenStudy (anonymous):

Help with this problem? Two end points of a line segment are (-8, 0) and (4, -2). What are the coordinates of the point on the line through which its bisector passes?

OpenStudy (anonymous):



OpenStudy (anonymous):

Find x,y using Section formula....) hope that helps

OpenStudy (anonymous):

@cmp411 ....

hartnn (hartnn):

or you could simply use mid-point formula \(x= (x_1+x_2)/2 \\ y=(y_1+y_2)/2\) (x1,y1)=(-8,0) (x2,y2)=(4,-2)